# Collaborative Unlabeled Data Optimization

### v1

@@ Abstract
This paper pioneers a novel data-centric paradigm to maximize the utility of unlabeled data, tackling a critical question: How can we enhance the efficiency and sustainability of deep learning training by optimizing the data itself? We begin by identifying three key limitations in existing model-centric approaches, all rooted in a shared bottleneck: knowledge extracted from data is locked to model parameters, hindering its reusability and scalability. To this end, we propose CoOpt, a highly efficient, parallelized framework for collaborative unlabeled data optimization, thereby effectively encoding knowledge into the data itself. By distributing unlabeled data and leveraging publicly available task-agnostic models, CoOpt facilitates scalable, reusable, and sustainable training pipelines. Extensive experiments across diverse datasets and architectures demonstrate its efficacy and efficiency, achieving 13.6% and 6.8% improvements on Tiny-ImageNet and ImageNet-1K, respectively, with training speedups of 1.94×1.94$\times$1.94 × and 1.2×1.2$\times$1.2 ×  [^1] Our code is provided in the Supplementary Materials and will be publicly accessible..
@@ 1 Introduction
Deep Learning has achieved remarkable success, primarily due to the large-scale datasets [28, 35]. Despite the abundance of data in the era of big data, a significant portion of them remains unlabeled [22]. The dominant paradigm in the field for exploiting unlabeled data is self-supervised learning (SSL), which is fundamentally model-centric: it carefully crafted pretext tasks and loss functions to encode data information into model parameters [6, 13, 4].
@@
However, the model-centric nature leads to three fundamental challenges. First, their training protocols are tightly coupled to specific network architectures, significantly hindering the transferability of trained model to other architectures [31, 18]. Second, even with advances in acceleration techniques, training over extensive unlabeled datasets still incurs prohibitively high computational costs [29]. Third, when a new downstream task arises, pretrained models typically require extensive fine-tuning or even full retraining. At the core of these challenges is a shared bottleneck: knowledge extracted from data is locked in model parameters, restricting its adaptability and preventing efficient reuse across diverse tasks or architectures.
@@
To break free from model-centric paradigm, we propose a data-centric paradigm that directly optimizes the unlabeled data by optimizing targets for samples (further elaborated in Section 3.1), thereby effectively encoding knowledge into the data itself rather than into model parameters. The resulting “optimal data” is agnostic to downstream architectures, accelerates subsequent training by providing richer supervision, and can be reused across multiple tasks without repeated large-scale pretraining.
@@
In the meanwhile, scaling this approach to massive unlabeled datasets introduces a significant challenge: a single node faces prohibitive compute and storage demands. To address this, we propose CoOpt, a highly efficient collaborative framework inspired by crowd-sourcing to achieve parallel data optimization. An overview of CoOpt is depicted in Figure 1, with detailed processes shown in Figure 2. In CoOpt, the unlabeled dataset is partitioned into disjoint subsets, each processed independently and in parallel by participants equipped with task-agnostic models. These models, referred to as prior models, can diverge from publicly available pre-trained models and the participants’ local models. Once the targets are optimized, they are aggregated to reconstruct a fully optimized dataset, achieving computational scalability through decentralized workload distribution.
CoOpt offers three key advantages: First, as the optimization relies solely on task-agnostic models, the aggregated optimized targets can be directly applied to any downstream architecture. Second, by distributing data subsets across participants, each node incurs only a fraction of the total computational load, enabling scalability and resource efficiency. Importantly, since only targets $D_{Y}$ are optimized, the aggregated targets are the same as those obtained by optimizing the whole dataset. Third, downstream tasks can directly fine-tune the optimized dataset for task-specific objectives without requiring training from scratch. Comprehensive experimental results have demonstrated that CoOpt achieves superior performance, generalization, and training efficiency.
@@
    We propose CoOpt, the first data-centric framework for optimizing unlabeled data, leveraging distributed computation to significantly lower the computational and storage costs per participant.
@@
    Within CoOpt, we identify a critical issue: Target Distribution Inconsistency in Section 3.3. Accordingly, we propose an effective target alignment strategy, elaborated in Section 3.4.
@@
    Extensive experiments have been conducted across a range of real-world scenarios, involving a variety of prior models trained on various datasets, architectures, and training paradigms. Notably, special cases are explored where human or significantly weak models are employed as prior models, verifying the robustness and flexibility of CoOpt. These experiments consistently demonstrate that CoOpt achieves superior effectiveness and efficiency.
@@ 2 Related Work
Model-centric Perspective: Self-supervised Learning.
It aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc [34] uses instance discrimination as a pretext task. MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR [6] highlights the importance of hard positive sample strategies. Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
Data-centric Perspective: Knowledge Distillation and Dataset Distillation.
Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11]. Many following works [45, 44, 11] aim to enhance the use of soft labels. Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into three primary approaches [22]: meta-learning frameworks [33, 46], matching-based methods [43, 42, 14] and decoupling frameworks [38, 29]. A more detailed description of knowledge distillation and dataset distillation is provided in Section A.
@@ 3 Collaborative Data Optimization Framework CoOpt
We begin by formally defining data optimization in Section 3.1. Subsequently, we provide a detailed description of the proposed CoOpt in Section 3.2. Furthermore, we identify an inherent challenge within this framework in Section 3.3 and present method in Section 3.4.
Table 1: Properties of Various Data Utilization Methods. ‘Optimize $D_{X}$’ and ‘Optimize $D_{Y}$’ indicate whether the optimization process encompasses the samples and targets, respectively. ‘SSL’ is self-supervised learning, ‘KD’ is knowledge distillation, and ‘DD’ is dataset distillation. ‘$\mathbf{c}$’ denotes the training cost of standard supervised learning that employs human-labeled data.
Method Optimize Efficiency Cost Analysis Total Efficiency
$D_{X}$ $D_{Y}$
Data
Optimization
Model
Training
Original SSL 0 [Uncaptioned image] ($\geq 2\mathbf{c}$) Extensive Computation [Uncaptioned image]
Labeled Data Optimization KD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Teacher Models [Uncaptioned image]
DD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Pre-trained models [Uncaptioned image]
CoOpt (Our) [Uncaptioned image] ($<\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Various Prior Models [Uncaptioned image]
@@ 3.1 Definition of Data Optimization
We first revisit current training acceleration techniques from a data-centric perspective, including knowledge distillation [17] and dataset distillation [33]. Specifically, we decouple data into samples $D_{X}$ and targets $D_{Y}$  [^2] A formal definition of $D_{X}$ and $D_{Y}$ is provided in Section 3.1., as summarized in Table 1. Essentially, compared to the self-supervised learning methods, these approaches achieve higher efficiency by either optimizing targets $D_{Y}$ pre-trained models (as in knowledge distillation), or by optimizing both input data $D_{X}$ and targets $D_{Y}$ (as in dataset distillation). Notably, for dataset distillation, optimizing input data $D_{X}$ is computationally more expensive than optimizing targets $D_{Y}$ [2]. Furthermore, recent studies [27, 25] have indicated that solely optimizing $D_{Y}$ not only reduces computational overhead but also achieves significant performance gains. These findings have demonstrated that optimizing $D_{Y}$ is both necessary and efficient. We refer to this process as data optimization.
@@
Notably, directly extending these methods to unlabeled data presents two challenges. First, they rely on task-specific models trained on labeled data, which are unavailable. Second, they incur substantial costs. For example, most dataset distillation methods rely on bi-level optimization [23]. To address these limitations, we propose a collaborative framework that leverages distributed computation and various task-agnostic models for unlabeled data. Specifically, inspired by [29], which shows that using task-agnostic models for target assignment can expedite training, we further enhance efficiency by splitting the data and then parallelly optimizing each split. After optimization, the optimal subset are aggregated to reconstruct a fully optimized dataset, achieving computational efficiency. Formally, we define data optimization with a prior model $\boldsymbol{\psi}$ in each participant in Proposition 1. {propositionframe}
Proposition 1 (Data optimization with prior model $\boldsymbol{\psi}$) .
@@
where $\mathbf{y}_{i}$ is the optimized target, and $\boldsymbol{\psi}(\mathbf{x}_{i})$ represents the target of $\mathbf{x}_{i}$, which means the feature representation. $\mathbf{W}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ denotes a matrix designed to transform the feature vector $\boldsymbol{\psi}(\mathbf{x}_{i})$ from dimension $l$ to $n$ without loss of information [24]. This transformation aligns the output dimension [^3] Here, $n$ is the target dimensionality of $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$. In practice, each participant produces targets of varying dimensions due to using different prior models. Therefore, to train the model $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$ on the optimized data $D^{\prime}$, we employ a random matrix $\mathbf{W}$ to transform all target vectors to a common dimensionality. with that required by the model trained on optimized data $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}:\mathbb{R}^{m}\to\mathbb{R}^{n}$.
@@ 3.2 Overview of the Proposed Framework CoOpt
CoOpt is a collaborative and parallelized framework that comprises an open data platform and $K$ participants, each equipped with a distinct prior model. Specifically, CoOpt operates through the following five steps:
@@
Step 2: Data optimization. Each participant optimizes their respective unlabeled dataset $D^{(k)}$ using the prior model $\boldsymbol{\psi}^{k}$. This data optimization process, defined in Proposition 1, yields optimized targets $D^{(k)}_{Y}$ and optimized data ${D^{\prime}}^{(k)}$.
Step 3: Data alignment. The heterogeneity of prior models among participants induces significant variations in their distribution of the targets. This issue, referred to as target distribution inconsistency (defined in Section 3.3), necessitates an alignment strategy (detailed in Section 3.4) to align the target distribution spaces across participants. Crucially, each participant needs to align their targets to the distribution space of the most optimal prior model using a learnable transformation matrix $\mathbf{T}^{(k)}$, yielding the final optimized dataset ${D^{\prime}}^{(k)}$. Further details are provided in Section 3.4.
@@
The proposed collaborative parallel process enables participants to independently and parallelly optimize their subsets while ensuring consistency through the proposed alignment strategies. Consequently, this approach markedly reduces individual data optimization costs and enhances data processing efficiency through parallel execution.
@@ 3.3 An Inherent Challenge: Target Distribution Inconsistency
In our collaborative framework, each participant may employ a distinct prior model, leading to inconsistencies in the target distributions, as illustrated in Figure 5(a). For example, participant 1 uses ResNet-18 for optimization, resulting in a target dimension of $512$, while participant 2 utilizes ResNet-50, yielding a target dimension of $2{,}048$. Such inconsistencies can negatively impact the generalization capabilities of models trained on the optimized data, as they prevent the models from learning representations that are uniformly representative of the overall data distribution.
@@ 3.4 An Effective Strategy: Target Alignment
To address this issue, a potential solution is to align the target distributions of all participants’ prior models with that of the prior model producing the most optimal target distribution space, referred to as the best prior model. Such alignment can be achieved by utilizing an optimizable transformation matrix to map each participant’s target distribution to that of the best prior model [29]. This alignment strategy ensures consistency across all optimized target distribution spaces.
@@
Drawing inspirations from [32], which proposes an optimizable metric a.k.a. uniform value loss to achieve feature uniformity on the hypersphere during training, we employ this metric to evaluate the quality of prior models. Notably, [32] also provide theoretical validation of the connection between uniformity and feature quality. Specifically, each participant downloads a small shared dataset $S_{X}$ from the platform and computes the uniformity value of their prior model on $S_{X}$. They then upload this value to the platform, enabling it to determine which participant possesses the best prior model. The uniform value is computed as:
@@
where $\boldsymbol{\psi}$ is the prior model, $\tau$ is a hyper-parameter set as $2$, consistent with [32].
A lower uniform value indicates a higher-quality prior model, which optimizes targets of superior quality. Extensive experiments in Figure 3(c) demonstrate a strong correlation between this metric and the performance of prior, thereby effectively assessing the quality of targets.
@@
Upon identifying the best prior model, denoted as $\boldsymbol{\psi}^{\star}$, all participants, excluding the best prior model itself, proceed to train an optimizable transformation matrix. Specifically, the participant owning the best prior model computes its optimized targets, denoted as ${\mathbf{S}_{Y}}^{\star}$, onthe shared dataset $S_{X}$ using the prior model ${\boldsymbol{\psi}}^{\star}$. ${\mathbf{S}_{Y}}^{\star}$ are then uploaded to the platform, which ensures they are publicly accessible to all participants. Following this, each participant $k$ optimizes a lightweight transformation matrix, denoted as $\mathbf{T}^{(k)}$, on the shared dataset $S_{X}$. The optimization problem is defined as follows:
@@
where $\mathbf{S}_{X}$ represents the matrix form of $S_{X}$, suitable for input into the network ${\boldsymbol{\psi}}^{(k)}$, and ${\mathbf{S}_{Y}}^{\star}$ also represents the matrix form of ${S_{Y}}^{\star}$. After obtaining the transformation matrix $\mathbf{T}^{(k)}$, the participant can convert the optimized targets for its own data using this matrix: ${D_{Y}}^{(k)}=\mathbf{T}^{(k)}\cdot{\boldsymbol{\psi}}^{(k)}({\mathbf{D}_{X}}^{(k)})$, where ${\mathbf{D}_{X}}^{(k)}$ denotes the participant’s subset, and ${D_{Y}}^{(k)}$ are the adjusted targets aligned with the best prior model’s target distribution space. As illustrated in Figure 5(b), the proposed alignment strategy effectively mitigates target distribution space inconsistency.
@@
Our proposed framework builds upon well-established theoretical foundations. Specifically, RELA [29] has theoretically demonstrated that leveraging task-agnostic models, such as pre-trained models, can accelerate learning. Notably, our framework diverges substantially from RELA in both its objectives and methodology. Specifically, CoOpt introduces a collaborative optimization approach tailored for unlabeled data, targeting a fundamentally distinct problem domain. The challenges associated with collaborative data optimization are unique to CoOpt and remain unaddressed by RELA. Furthermore, the metric we employ to evaluate the quality of prior models, uniform value loss [32], has been theoretically validated for its effectiveness.
Dataset Metric BYOL DINO MoCo SimCLR SimSiam SwAV DCL CoOpt
CF-10 Acc. (%) 82.8 $\pm$ 0.1 82.6 $\pm$ 0.0 82.9 $\pm$ 0.1 83.1 $\pm$ 0.0 79.0 $\pm$ 0.0 82.9 $\pm$ 0.1 83.9 $\pm$ 0.1 88.7 $\pm$ 0.1 ($\uparrow$ 4.8)
CF-100 Acc. (%) 51.7 $\pm$ 0.1 51.0 $\pm$ 0.0 57.8 $\pm$ 0.1 55.4 $\pm$ 0.0 44.6 $\pm$ 0.1 53.2 $\pm$ 0.1 58.2 $\pm$ 0.2 65.8 $\pm$ 0.1 ($\uparrow$ 7.6)
T-IN Acc. (%) 43.9 $\pm$ 0.2 36.1 $\pm$ 0.0 42.4 $\pm$ 0.2 41.5 $\pm$ 0.1 40.8 $\pm$ 0.0 39.9 $\pm$ 0.1 44.6 $\pm$ 0.0 58.2 $\pm$ 0.1 ($\uparrow$ 13.6)
IN-1K Acc.(%) 61.9 $\pm$ 0.1 52.2 $\pm$ 0.0 57.6 $\pm$ 0.0 58.0 $\pm$ 0.0 55.8 $\pm$ 0.1 57.2 $\pm$ 0.1 60.6 $\pm$ 0.1 68.7 $\pm$ 0.0 ($\uparrow$ 6.8)

### v2

# Beyond Model-centric: Collaborative DATA Optimization for Reusing and Sharing

@@ Abstract
This paper pioneers a novel data-centric paradigm to maximize the utility of unlabeled data, tackling a critical question: How can we enhance the sustainability and efficiency of deep learning training by optimizing the data itself? We begin by identifying two key limitations in existing model-centric approaches, all rooted in a shared bottleneck: knowledge extracted from data is locked to model parameters, hindering its reusability and scalability. To this end, we propose CoOpt, a highly efficient, parallelized framework for collaborative unlabeled data optimization. By distributing unlabeled data and leveraging publicly available task-agnostic prior models, CoOpt optimizes raw unlabeled data into knowledge-enriched training sets that are effective, efficient, reusable, and easily shareable. Extensive experiments across diverse datasets and architectures validate these advantages, achieving a 7.9% improvement on ImageNet-1K over BYOL. Notably, CoOpt remains effective even when all prior models are significantly weak, substantially accelerating the early stages of training. These results establish data-centric optimization as a promising path toward sustainable and efficient deep learning  [^1] Code will be publicly available..
@@ 1 Introduction
Deep Learning has achieved remarkable success, primarily due to the large-scale datasets (Song et al., 2020; Yang et al., 2023). Despite the abundance of data in the era of big data, a significant portion of them remains unlabeled (Lei & Tao, 2023). The dominant paradigm in the field for exploiting unlabeled data is self-supervised learning (SSL), which is fundamentally model-centric: it carefully crafted pretext tasks and loss functions to encode data information into model parameters (Chen et al., 2020a; Grill et al., 2020; Gui et al., 2024).
@@
However, the model-centric nature presents two critical challenges. First, their training protocols are tightly coupled to specific network architectures, severely hindering the transferability and reusability of trained model to other architectures (Wagner et al., 2022; Huang et al., 2023). Second, despite acceleration advances, training over extensive unlabeled datasets still computationally prohibitive (Sun et al., 2024). At the core of these challenges is a shared bottleneck: knowledge extracted from data is locked in model parameters, restricting its adaptability and preventing efficient reuse across diverse tasks or architectures.
@@
To break free from model-centric paradigm, we propose a data-centric paradigm that directly optimizes the unlabeled data by optimizing targets for samples (detailed in Sec. 3.1), thereby effectively encoding knowledge into the data itself rather than into model parameters. The resulting “optimal data” is agnostic to downstream architectures, accelerates subsequent training by providing richer supervision, and can be reused across multiple tasks without repeated large-scale pretraining.
@@
In the meanwhile, scaling this approach to massive unlabeled datasets introduces a significant challenge: a single node faces prohibitive compute and storage demands. To address this, we propose CoOpt, a highly efficient collaborative framework inspired by crowd-sourcing to achieve parallel data optimization. An overview of CoOpt is depicted in Fig. 1, with detailed processes shown in Fig. 3. In CoOpt, the unlabeled dataset is partitioned into disjoint subsets, each processed independently and in parallel by participants equipped with task-agnostic models. These models, referred to as prior models, can diverge from publicly available pre-trained models and the participants’ local models. Once the targets are optimized, they are aggregated to reconstruct a fully optimized dataset, achieving computational scalability through decentralized workload distribution.
CoOpt offers three key advantages: First, relying solely on task-agnostic prior models, the optimized data can be directly used to any downstream architecture, thereby ensuring strong generalizability and reusability (see Sec. 4.2). Second, by distributing non-overlapping data subsets across participants, each node handles only a fraction of the total computational cost, thereby enabling scalable and resource-efficient optimization (see Sec. 4.2). Third, CoOpt is lightweight, incurring negligible overhead while substantially enhancing training efficiency and performance (see Sec. 4.4).
@@
    We propose CoOpt, the first data-centric framework for collaboratively optimizing unlabeled data. By leveraging task-agnostic prior models, CoOpt transforms raw unlabeled samples into optimal data, enabling high performance, efficiency, strong generalization, and reusability.
@@
    Within CoOpt, we identify a critical issue, Target Distribution Inconsistency (Sec. 3.3), and introduce a lightweight target alignment strategy to address it (Sec. 3.4).
@@
    We conduct experiments across datasets and models to comprehensively validate the advantages of CoOpt (Sec. 4.2). Further, we provide a detailed analysis of the key factors influencing its effectiveness (Sec. 4.3). Remarkably, we demonstrate that CoOpt remains effective even when all prior models are weak, substantially accelerating the early stages of training.
@@ 2 Related Work
Self-Supervised Learning.
It aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc (Wu et al., 2018) uses instance discrimination as a pretext task. MoCo (He et al., 2020) significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR (Chen et al., 2020a) highlights the importance of hard positive sample strategies. Notably, BYOL (Grill et al., 2020) discards negative sampling and surpasses the performance of SimCLR (Chen et al., 2020a).
Model-Centric Perspective: Knowledge Distillation.
Knowledge distillation (Hinton, 2015) leverages teacher-generated soft labels to improve student training efficiency and performance (Dong et al., 2023). A line of knowledge distillation methods utilizes multiple teachers (MKD) (Zhang et al., 2022; Pham et al., 2023) to enhance student learning. They assume that ensemble outputs from multiple teachers enables students to learn more generalized representations. Notably, all teacher models process the same input data.
Our setting departs fundamentally from knowledge distillation in terms of objective, input data, and teacher models (see Fig. 2). First, in KD, the distilled knowledge is embedded in student parameters, limiting its reuse across different architectures. In contrast, our objective is to construct a high-quality, optimized dataset that is model-agnostic and reusable, enabling training or evaluation of diverse architectures. Second, rather than feeding all teachers the same inputs, we partition the unlabeled data into disjoint subsets, each optimized by a different prior model. Third, existing KD methods often rely on intricate loss functions (Jiang et al., 2024) or require teacher fine-tuning (Wu et al., 2021), but our framework uses arbitrary pre-trained models without domain-specific adaptation. The optimized data can then be directly reused to train arbitrary downstream models without further modification.
Data-Centric Perspective: Dataset Distillation.
Dataset distillation (Wang et al., 2018) improves the training efficiency by learning a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into three primary approaches (Lei & Tao, 2023): meta-learning frameworks (Wang et al., 2018; Zhou et al., 2022), matching-based methods (Zhao et al., 2020; Zhao & Bilen, 2023; Guo et al., 2024) and decoupling frameworks (Yin et al., 2023; Sun et al., 2024). Notably, current methods predominantly focus on distilling labeled datasets.
@@ 3 Collaborative Data Optimization Framework CoOpt
We begin by formally defining data optimization in Sec. 3.1. Subsequently, we provide a detailed description of the proposed CoOpt in Sec. 3.2. Furthermore, we identify an inherent challenge within this framework in Sec. 3.3 and present method in Sec. 3.4.
Figure 2: Comparison Between KD and Ours.
@@ 3.1 Definition of Data Optimization
We first revisit current training acceleration techniques, including knowledge distillation (Hinton, 2015) and dataset distillation (Wang et al., 2018). Specifically, we decouple data into samples $D_{X}$ and targets $D_{Y}$. Essentially, compared to self-supervised learning methods, these approaches are more efficient. They achieve this by either optimizing the target $D_{Y}$ of pre-trained models (as in knowledge distillation) or by jointly optimizing both the input data $D_{X}$ and target $D_{Y}$ (as in dataset distillation). Notably, for dataset distillation, optimizing input data $D_{X}$ is computationally more expensive than optimizing targets $D_{Y}$ (Bohdal et al., 2020). Furthermore, recent studies Shang et al. (2025); Qin et al. (2024) have indicated that solely optimizing $D_{Y}$ not only reduces computational overhead but also achieves significant performance gains. These findings have demonstrated that optimizing $D_{Y}$ is both necessary and efficient. We refer to this process as data optimization.
@@
Notably, directly extending these methods to unlabeled data is infeasible. Existing dataset distillation methods focus on labeled data, whereas knowledge distillation methods, as reviewed in Sec. 2, significantly differ from ours. A detailed comparison is presented in Fig. 2. To address these limitations, we propose a collaborative framework that leverages distributed computation and various task-agnostic models for unlabeled data. Specifically, inspired by (Sun et al., 2024), which shows that using task-agnostic models for target assignment can expedite training, we further enhance efficiency by splitting the data and then parallelly optimizing each split. After optimization, the optimal subset are aggregated to reconstruct a fully optimized dataset, achieving computational efficiency. Formally, we define data optimization with a prior model $\boldsymbol{\psi}$ in each participant in Def. 1. {definitionframe}
Definition 1 (Data optimization with prior model $\boldsymbol{\psi}$) .
@@
where $\mathbf{y}_{i}$ is the optimized target, and $\boldsymbol{\psi}(\mathbf{x}_{i})$ represents the target of $\mathbf{x}_{i}$, which means the feature representation. $\mathbf{W}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ denotes a matrix designed to transform the feature vector $\boldsymbol{\psi}(\mathbf{x}_{i})$ from dimension $l$ to $n$ without loss of information (Matoušek, 2008). This transformation aligns the output dimension [^2] Here, $n$ is the target dimensionality of $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$. In practice, each participant produces targets of varying dimensions due to using different prior models. Therefore, to train the model $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$ on the optimized data $D^{\prime}$, we employ a random matrix $\mathbf{W}$ to transform all target vectors to a common dimensionality. with that required by the model trained on optimized data $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}:\mathbb{R}^{m}\to\mathbb{R}^{n}$.
@@ 3.2 Overview of the Proposed Framework CoOpt
CoOpt is a collaborative and parallelized framework that includes an open data platform and $K$ participants, each equipped with a distinct prior model. CoOpt operates through the five steps:
@@
Step 2: Data Optimization. Each participant optimizes their respective unlabeled dataset $D^{(k)}$ using the prior model $\boldsymbol{\psi}^{k}$. This data optimization process, defined in Prop. 1, yields optimized targets $D^{(k)}_{Y}$ and optimized data ${D^{\prime}}^{(k)}$.
Step 3: Data Alignment. The heterogeneity of prior models among participants induces significant variations in their distribution of the targets. This issue, referred to as target distribution inconsistency (defined in Sec. 3.3), necessitates an alignment strategy (detailed in Sec. 3.4) to align the target distribution across participants. Crucially, each participant needs to align their targets distribution to the most optimal prior model using a learnable transformation matrix $\mathbf{T}^{(k)}$, yielding the final optimized dataset ${D^{\prime}}^{(k)}$. Further details are provided in Sec. 3.4.
@@
The proposed CoOpt enables participants to independently and parallelly optimize their subsets while ensuring consistency through the proposed alignment strategies. Consequently, this approach markedly reduces individual data optimization costs and enhances efficiency.
@@ 3.3 An Inherent Challenge: Target Distribution Inconsistency
In our collaborative framework, each participant may employ a distinct prior model, leading to inconsistencies in the target distributions, as illustrated in Fig. 6(a). For example, participant 1 uses ResNet-18 for optimization, resulting in a target dimension of $512$, while participant 2 utilizes ResNet-50, yielding a target dimension of 2,0482{,}048. Such inconsistencies can negatively impact the generalization capabilities of models trained on the optimized data, as they prevent the models from learning representations that are uniformly representative of the overall data distribution.
@@ 3.4 An Effective Strategy: Target Alignment
To address this issue, a potential solution is to align the target distributions of all participants’ prior models with that of the prior model producing the most optimal target distribution, referred to as the best prior model. Such alignment can be achieved by utilizing an optimizable transformation matrix to map each participant’s target distribution to that of the best prior model (Sun et al., 2024). This alignment strategy ensures consistency across all optimized target distributions.
@@
Inspired by Wang & Isola (2020), which proposes an optimizable metric a.k.a. uniform value loss to achieve feature uniformity on the hypersphere during training, we employ this metric to evaluate the quality of prior models. Notably, (Wang & Isola, 2020) also provide theoretical validation of the connection between uniformity and feature quality. Specifically, each participant downloads a small shared dataset $S_{X}$ from the platform and computes the uniformity value of their prior model on $S_{X}$. They then upload this value to the platform, enabling it to determine which participant possesses the best prior model. The uniform value is computed as:
@@
where $\boldsymbol{\psi}$ is the prior model, $\tau$ is a hyper-parameter set as $2$, consistent with (Wang & Isola, 2020).
A lower uniform value indicates a higher-quality prior model, which optimizes targets of superior quality. Extensive experiments in Fig. 5(c) demonstrate a strong correlation between this metric and the performance of prior, thereby effectively assessing the quality of targets.
@@
Upon identifying the best prior model, denoted as $\boldsymbol{\psi}^{\star}$, all participants, excluding the best prior model itself, proceed to train an optimizable transformation matrix. Specifically, the participant owning $\boldsymbol{\psi}^{\star}$ computes its optimized targets, denoted as $\mathbf{S}_{Y}$$^{\star}$, on the shared dataset $S_{X}$. $\mathbf{S}_{Y}$$^{\star}$ are then uploaded to the platform, which ensures they are publicly accessible to all participants. Following this, each participant $k$ optimizes a lightweight transformation matrix, denoted as $\mathbf{T}^{(k)}$, on the shared dataset $S_{X}$. The optimization problem is defined as follows:
@@
where $\mathbf{S}_{X}$ represents the matrix form of $S_{X}$, suitable for input into the network ${\boldsymbol{\psi}}^{(k)}$, and $\mathbf{S}_{Y}$$^{\star}$ also represents the matrix form of $S_{Y}$$^{\star}$. After obtaining the transformation matrix $\mathbf{T}^{(k)}$, the participant can convert the optimized targets for its own data using this matrix: ${D_{Y}}^{(k)}=\mathbf{T}^{(k)}\cdot{\boldsymbol{\psi}}^{(k)}({\mathbf{D}_{X}}^{(k)})$, where ${\mathbf{D}_{X}}^{(k)}$ denotes the participant’s subset, and ${D_{Y}}^{(k)}$ are the adjusted targets aligned with the best prior model’s target distribution. As illustrated in Fig. 6(b), the proposed alignment strategy effectively mitigates target distribution inconsistency.
@@
CoOpt builds upon well-established theoretical foundations. In particular, RELA Sun et al. (2024) has theoretically demonstrated that leveraging task-agnostic models, such as pre-trained models, can accelerate model learning. Nevertheless, CoOpt departs significantly from RELA in both its research objectives and methodological design. Specifically, CoOpt introduces a collaborative optimization approach tailored for unlabeled data, targeting a fundamentally distinct problem domain. The challenges associated with collaborative data optimization are unique to CoOpt and remain unaddressed by RELA. Furthermore, the metric we employ to evaluate the quality of prior models, uniform value loss (Wang & Isola, 2020), has been theoretically validated for its effectiveness.